OpenStudy (anonymous):

What is the y-intercept of -10x-1/2y=5

OpenStudy (primeralph):

Plug in x = 0.

OpenStudy (anonymous):

so, -10(0)-1/2y=5?

OpenStudy (primeralph):

Yeah, solve for y.

OpenStudy (anonymous):

is it -2?

OpenStudy (primeralph):

No. How did you get that?

OpenStudy (anonymous):

actually -2.5 sorry. 0-1/2y=5 -1/2y=5 y=-2.5

OpenStudy (primeralph):

Nope. Small error.

OpenStudy (anonymous):

whats that?

OpenStudy (primeralph):

Last line.

OpenStudy (anonymous):

so is it -1/2 ?

OpenStudy (primeralph):

No.

OpenStudy (primeralph):